### v1

# Interpretability-Guided Bi-objective Optimization: Aligning Accuracy and Explainability

Abstract
This paper introduces Interpretability-Guided Bi-objective Optimization (IGBO), a framework that trains interpretable models by incorporating structured domain knowledge via a bi-objective formulation. IGBO encodes feature importance hierarchies as a Directed Acyclic Graph (DAG) and uses Temporal Integrated Gradients (TIG) to measure feature importance. To address the Out-of-Distribution (OOD) problem in TIG computation, we propose an Optimal Path Oracle that learns data-manifold-aware integration paths. Theoretical analysis proves convergence properties and robustness to mini-batch noise, while empirical results on time-series data demonstrate IGBO’s effectiveness in enforcing DAG constraints with minimal accuracy loss, outperforming standard regularization baselines.
Keywords: Interpretable Machine Learning, Bi-objective Optimization, Explainable AI (XAI), Temporal Integrated Gradients, Optimal Path Oracle, Structured Regularization
Note: AI writing assistance was utilized for improved readability while preserving all technical integrity.
I Introduction
Deep learning models have achieved remarkable success in sequential prediction tasks across domains such as healthcare and finance. However, their inherent complexity often renders them opaque “black boxes,” limiting trust and deployment in safety-critical scenarios where understanding the model’s reasoning is as important as its accuracy. Post-hoc explanation methods, including Integrated Gradients [sundararajan2017axiomatic] and LIME [ribeiro2016should], attempt to illuminate this reasoning by analyzing trained models. A fundamental limitation persists: these methods are applied after training, providing no guarantee that the model’s internal decision-making aligns with established domain knowledge or desired behavioral constraints during the learning process itself [mironicolau2024comprehensive]. Consequently, a model may achieve high accuracy by exploiting spurious, non-causal, or ethically problematic correlations in the data—a risk unacceptable in high-stakes applications.
To build more trustworthy systems, interpretability must be integrated directly into the training objective. A common approach adds a regularization term penalizing undesirable behavior. However, such methods often treat interpretability as a monolithic, soft constraint, offering practitioners limited control over the precise trade-off between accuracy and interpretability. More critically, they lack a formal mechanism to enforce structured, relational constraints that reflect complex domain expertise, such as hierarchies of feature importance. Additionally, gradient-based attribution methods like Integrated Gradients suffer from the Out-of-Distribution (OOD) problem when computing gradients along straight-line paths between inputs and baselines.
This work proposes a principled framework to address these gaps: Interpretability-Guided Bi-objective Optimization (IGBO). IGBO formalizes interpretable model training as two interconnected bi-objective optimization problems: (1) training the main model $F_{\theta}$ to simultaneously minimize a primary task loss $\mathcal{L}(\theta)$ and an interpretability loss $\mathcal{H}(\theta)$, and (2) training an Optimal Path Oracle $G$ to generate data-manifold-aware integration paths for robust gradient computation. The interpretability loss is derived from two key components: (1) a Temporal Integrated Gradients (TIG) metric that provides a differentiable measure of feature importance over time, and (2) a Directed Acyclic Graph (DAG) defined by a domain expert, which encodes permissible relative importance relationships between features (e.g., feature $A$ should be more important than feature $B$).
Optimal Path Oracle: To address the OOD problem in TIG computation, we introduce a learnable oracle that generates sequences of intermediate points forming the shortest possible path within the data distribution. This oracle is trained with dual objectives: minimizing path length and ensuring generated points are in-distribution. During integration, we employ a flexible strategy using $K$ oracle-generated anchor points and MKM-K interpolated points, creating a tunable computational trade-off between path fidelity and efficiency.
Gradient-Based Optimization: Both bi-objective problems require navigating trade-offs between competing gradients. We employ a geometric gradient combination rule based on the cosine similarity between gradient vectors. When gradients are aligned, we use a convex combination; when they conflict, we combine their orthogonal components. This approach provides descent guarantees for both objectives while avoiding the need for secondary optimization subroutines.
Contributions. This paper makes the following contributions:
  • We introduce the IGBO framework, which integrates structured domain knowledge (via a DAG) into gradient-based learning through a differentiable TIG-based importance function $H_{k}(\mathbf{X},\theta)$, addressing both interpretability constraints and OOD issues in attribution.
  • We propose an Optimal Path Oracle that learns to generate data-manifold-aware integration paths, enabling robust TIG computation while providing a flexible trade-off between path quality and computational cost via parameters $K$ and $M$.
  • We establish a rigorous empirical validation protocol to assess IGBO’s capability to enforce complex DAG constraints on both synthetic and real-world time-series data, demonstrating its effectiveness with minimal compromise on predictive accuracy.
II Related Work
Our work synthesizes concepts from interpretable AI, gradient-based multi-objective optimization, robust attribution methods, and the geometry of loss landscapes. We position IGBO within these interconnected research streams.
Integrating Interpretability into Training: Moving beyond post-hoc explanation methods like LIME [ribeiro2016should] and SHAP [lundberg2017unified], recent approaches integrate interpretability directly into the training loop. Methods like Right for the Right Reasons [ross2017right] regularize input gradients to align with human annotations, focusing on local, per-instance explanations. Concept Bottleneck Models (CBMs) [koh2020concept] enforce interpretability by routing predictions through human-specified concepts, requiring concept-level labels. Other works use symbolic constraints or logic rules to guide learning [marra2019lyra]. In contrast, IGBO enforces global, relational constraints on aggregate feature importance via a domain-expert DAG, without requiring per-instance annotations or explicit concept labels. This provides a structured prior over model behavior, complementing approaches that integrate domain knowledge to ensure explanation fidelity [mironicolau2024comprehensive, rudin2019stop]. Our work also relates to methods imposing sparsity or simplicity constraints but differs in its focus on explicit, relational importance hierarchies.
Multi-Objective Gradient-Based Optimization: IGBO relies on gradient-based techniques for multi-objective optimization. The Multi-Gradient Descent Algorithm (MGDA) [desideri2012multiple, fliege2019steepest] provides a general framework for finding descent directions common to multiple objectives, often requiring solving a quadratic program. For the prevalent two-objective case, IGBO’s geometric update rule provides a closed-form solution equivalent to MGDA, eliminating the optimization subroutine. This connects IGBO to gradient-based multi-task learning [sener2018multi], where conflicting gradients are a central challenge, and to Pareto-frontier tracing methods [lin2020pareto]. Recent work on multi-objective optimization includes adaptive weighting schemes [navon2022multi] and gradient manipulation techniques [liu2021conflict], but these typically do not consider interpretability-specific constraints as structured objectives.
Gradient-based Attribution and Robust Path Methods: Our Temporal Integrated Gradients (TIG) builds directly upon Integrated Gradients (IG) [sundararajan2017axiomatic], inheriting its axiomatic properties. Recent work has identified limitations in IG, particularly its sensitivity to the choice of baseline [sturmfels2020visualizing] and the fact that straight-line integration paths often traverse Out-of-Distribution (OOD) regions, leading to unreliable attributions [hess2020middle]. Several approaches attempt to address these issues: Expected Gradients [erion2019learning] averages over multiple baselines; BlurIG [xu2020explainable] applies smoothing; and adversarial training methods aim to produce more robust attributions [giove2023robust]. Our Optimal Path Oracle introduces a novel direction by learning a data-manifold-aware integration path, fundamentally changing how the attribution integral is computed. This approach is conceptually related to works exploring shortest paths on manifolds for explanations [garreau2020explaining] but differs in its fully differentiable, learned formulation and integration into a bi-objective training loop.
Geometry of Loss Landscapes and Gradient Alignment: Understanding and leveraging the geometry of high-dimensional loss landscapes is crucial for optimization [kiselev2024unraveling]. The geometric update rule in IGBO explicitly operates within the low-dimensional subspace spanned by two gradients. This connects to research on gradient conflict analysis in multi-task learning [yu2020gradient], where gradient directions are manipulated to improve convergence. The use of orthogonal gradient components when objectives conflict relates to techniques for escaping saddle points [dauphin2014identifying] and methods that project gradients to find mutually beneficial directions [schafer2021continuous]. IGBO’s approach is tailored specifically for navigating the accuracy-interpretability trade-off with clear geometric intuition, providing a practical instance of these geometric principles.
Evaluation of Explanations and Benchmarking: Assessing explanation quality remains a core challenge in XAI due to the frequent lack of ground truth [mironicolau2024comprehensive]. Recent work proposes synthetic benchmarks [agarwal2023evaluating], axiomatic evaluation frameworks [hedstrom2022quantus], and methods to measure explanation robustness [yeh2019completeness]. Our use of an expert-defined DAG provides a concrete, testable prior for evaluating the alignment of learned feature importance with domain knowledge, contributing to research on measurable explanation fidelity. This approach aligns with recent calls for more rigorous, application-grounded evaluation of interpretability methods [doshi2017towards, lyon2022human].
III Methodology
The Interpretability-Guided Bi-objective Optimization (IGBO) framework consists of two interconnected bi-objective optimization processes, both of which rely on gradient-based techniques for balancing competing objectives. These processes are:
  1. 1.
    Training the main sequential model $F_{\theta}$ to balance predictive accuracy and interpretability constraints;
  2. 2.
    Training an optimal path generator $G_{K}$ to produce data-manifold-aware integration paths for robust gradient computation in Temporal Integrated Gradients (TIG).
This section formalizes the interpretability prior, introduces the TIG-based importance measure, and details the proposed Routing Oracle for reliable path integration.
III-A Interpretability Prior Definition
Sequential Model Definition.
Let $\mathcal{X}=\mathbb{R}^{T\times d}$ denote the input space of multivariate time series of length $T$ and feature dimension $d$. A sequential model processes an input $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{T}]\in\mathcal{X}$ through a parameterized transition function $g_{\theta}$ and an output function $f_{\theta}$. The hidden state $\mathbf{h}_{t}\in\mathbb{R}^{h}$ at time $t$ evolves as:
$\mathbf{h}_{t}=g_{\theta}(\mathbf{x}_{t},\mathbf{h}_{t-1}),\quad t=1,\dots,T,$ (1)
with $\mathbf{h}_{0}=\mathbf{0}$. The output at time $t$ is given by $y_{t}=f_{\theta}(\mathbf{x}_{t},\mathbf{h}_{t-1})\in\mathbb{R}$. Consequently, the full sequence-to-sequence model $F_{\theta}:\mathcal{X}\to\mathbb{R}^{T}$ maps an input sequence to an output sequence, where its $t$-th component is $F_{\theta}(\mathbf{X})]_{t}=f_{\theta}(\mathbf{x}_{t},\mathbf{h}_{t-1})$.
Temporal Integrated Gradients (TIG).
To obtain a differentiable measure of feature importance over time, we extend Integrated Gradients [sundararajan2017axiomatic] to sequential data. For a given input $\mathbf{X}$ and a baseline $\mathbf{X}^{\prime}$, the importance of feature $k$ at time $t$ along a path $\gamma:[0,1]\to\mathcal{A}$ connecting $\gamma(0)=\mathbf{X}^{\prime}$ to $\gamma(1)=\mathbf{X}$ is defined as:
$\text{TIG}_{t,k}^{\gamma}(\mathbf{X},\theta)=\int_{0}^{1}\frac{\partial F_{\theta}(\gamma(\alpha))}{\partial\mathbf{X}_{t,k}}\cdot\frac{d\gamma_{t,k}(\alpha)}{d\alpha}\,d\alpha.$ (2)
For the special case of a straight-line path $\gamma_{\text{lin}}(\alpha)=\alpha\mathbf{X}+(1-\alpha)\mathbf{X}^{\prime}$, this simplifies to the standard IG formulation:
$\text{TIG}_{t,k}^{\gamma_{\text{lin}}}(\mathbf{X},\theta)=(\mathbf{X}_{t,k}-\mathbf{X}^{\prime}_{t,k})\int_{0}^{1}\frac{\partial F_{\theta}(\gamma_{\text{lin}}(\alpha))}{\partial\mathbf{X}_{t,k}}\,d\alpha.$ (3)
This integral aggregates the sensitivity of the model’s output to feature $(t,k)$ along the entire path $\gamma$. The standard implementation uses the straight‑line path $\gamma_{\text{lin}}(\alpha)=\alpha\mathbf{X}+(1-\alpha)\mathbf{X}^{\prime}$. However, a major drawback of this linear interpolation is that intermediate points $\gamma_{\text{lin}}(\alpha)$ often lie outside the training data distribution (Out‑of‑Distribution, OOD). Since $F_{\theta}$ is not trained to behave consistently in OOD regions, the resulting gradients—and hence the attributions—can be unreliable and noisy.
III-A1 Routing Oracle: Learning Data‑Manifold‑Aware Integration Paths
Motivation and Design Philosophy.
To circumvent the OOD problem inherent in straight‑line integration paths, we propose a Routing Oracle (also referred to as a Pathfinder) that learns to generate piecewise‑linear paths adhering to the data manifold while minimizing total path length. By ensuring all intermediate integration points reside within regions where the model has been properly trained, the oracle yields stable and reliable gradient computations for TIG.
Flexible Implementation Options.
The oracle is designed with practical deployment in mind, offering several configuration choices:
  • Baseline selection: The baseline $\mathbf{X}^{\prime}$ can be chosen per sample (e.g., using dataset statistics) or set globally (e.g., a zero sequence). For maximum simplicity, $\mathbf{X}^{\prime}$ can be fixed as a hyperparameter.
  • Path complexity control: The number of anchor points $K$ can be tuned according to the available computational budget. Larger $K$ allows the path to better follow the data manifold but increases inference cost. For deterministic high‑throughput deployment, $K$ can also be fixed as a hyperparameter.
This flexibility lets practitioners balance attribution quality against computational cost for their specific application.
Oracle Architecture and Recurrence Structure.
The routing oracle is implemented as a recurrent neural network $G:\mathcal{X}\times\mathcal{X}\times\mathbb{N}\to\mathcal{X}^{*}$ that, given a baseline $\mathbf{X}^{\prime}$, a target input $\mathbf{X}$, and a desired number of steps $K$, generates a sequence of $K$ anchor points:
$\mathbf{p}_{1},\dots,\mathbf{p}_{K}]=G(\mathbf{X}^{\prime},\mathbf{X},K;\theta_{G}).$
The architecture employs a recurrence that explicitly accounts for the remaining path length. Let $\mathbf{h}_{i}\in\mathbb{R}^{h}$ be a hidden state. Starting with $\mathbf{h}_{0}=\mathbf{0}$ and $\mathbf{p}_{0}=\mathbf{X}^{\prime}$, the $i$-th point is produced as:
$\mathbf{p}_{i}=r^{\text{output}}_{\theta_{G}}\bigl(K-i+1,\;\mathbf{X},\;\mathbf{p}_{i-1},\;\mathbf{h}_{i-1}\bigr),$
$\mathbf{h}_{i}=r^{\text{hidden}}_{\theta_{G}}\bigl(K-i+1,\;\mathbf{X},\;\mathbf{p}_{i-1},\;\mathbf{h}_{i-1}\bigr),$
where $r$ is a learned function. The explicit inclusion of the remaining‑step count (Ki+1)(K-i+1) allows the network to plan the path efficiently and adapt its behavior as the endpoint approaches.
Two‑Phase Training Procedure.
Training is performed in two distinct phases to ensure stability and reuse existing knowledge of the data distribution:
  1. 1.
    Discriminator pre‑training: We employ the discriminator $D(\cdot;\theta_{D})$ from a pre‑trained Generative Adversarial Network (GAN) as a frozen validity assessor. This discriminator, already trained to distinguish real data from synthetic samples, provides a reliable estimate of $\Pr(\mathbf{p}\in\mathcal{A})$ without requiring additional training.
  2. 2.
    Oracle training: With $D$ fixed, we train $G$ to minimize a composite loss that combines path shortness and point validity:
    $\mathcal{L}_{G}(\theta_{G})=\mathbb{E}_{(K,\mathbf{X}^{\prime},\mathbf{X})}\bigl[\mathcal{L}_{\text{path}}+\beta\cdot\mathcal{L}_{\text{valid}}\bigr],$ (4)
    where
    $\begin{split}&\mathcal{L}_{\text{path}}=\\ &\frac{\|\mathbf{X}^{\prime}-\mathbf{p}_{1}\|^{2}+\|\mathbf{X}-\mathbf{p}_{K}\|^{2}+\sum_{i=1}^{K-1}\|\mathbf{p}_{i+1}-\mathbf{p}_{i}\|^{2}}{K+1},\end{split}$ (5)
    $\mathcal{L}_{\text{valid}}=\frac{-1}{K}\sum_{i=1}^{K}\log D(\mathbf{p}_{i};\theta_{D}).$ (6)
    The hyper‑parameter $\beta>0$ controls the trade‑off between the two objectives.
Integration Strategy:
For the actual TIG computation with oracle-generated paths, we approximate the integral in Eq. (2) using numerical integration. Given $M$ integration points $\mathbf{q}_{1},\dots,\mathbf{q}_{M}$ along the path (including oracle anchors and interpolated points), we compute:
$\text{TIG}_{t,k}^{\text{Oracle}}(\mathbf{X},\theta)\approx\sum_{j=1}^{M-1}\frac{\partial F_{\theta}(\mathbf{q}_{j})}{\partial\mathbf{q}_{j,t,k}}\cdot(\mathbf{q}_{j+1,t,k}-\mathbf{q}_{j,t,k}),$ (7)
which corresponds to a Riemann sum approximation of the line integral along the piecewise-linear path connecting the oracle’s anchor points.
Transferability and Reusability.
A significant practical advantage of the Routing Oracle is its transferability across models. Once trained on a representative dataset $\mathcal{D}$, the same oracle can be applied to compute TIG for any model $F_{\theta}$ trained on $\mathcal{D}$, without retraining. This enables:
  • Consistent evaluation: All models are assessed using identical integration paths, allowing fair comparison of their feature‑importance patterns.
  • Computational economy: The costly training phase is performed only once per data domain.
  • Streamlined deployment: New models can be interpreted immediately using a pre‑existing, validated oracle, which is especially valuable in regulated domains (e.g., healthcare or finance) where explanation consistency is critical.
III-B Feature Importance Function $H_{k}(\mathbf{X},\theta)$
To obtain a single, global importance score for feature $k$ across the entire sequence for a specific input, we aggregate the temporal attributions. We define the importance function as:
$H_{k}(\mathbf{X},\theta)=\frac{\sum_{i=1}^{T}\sum_{j=1}^{T}\left|\text{TIG}_{i,k}(\mathbf{X},\theta)[j]\right|}{T(T+1)/2}.$ (8)
The denominator normalizes the score by the total absolute deviation of feature $k$ from its baseline, rendering $H_{k}$ a measure of the model’s sensitivity per unit change in the input feature. This normalization ensures importance scores are comparable across features with different native scales or units. The function $H_{k}:\mathcal{X}\times\Theta\to\mathbb{R}_{\geq 0}$ is central to our framework, providing a differentiable mapping from the joint space of inputs and parameters to a non-negative importance score.
Interpretability Graph (DAG) as a Structured Prior: Domain expertise is encoded as a Directed Acyclic Graph $G=(V,E)$. Each node $v\in V$ corresponds to an input feature. A directed edge $u\rightarrow v$ encodes the semantic prior that, on average across the data distribution $\mathcal{D}$, feature $u$ should be more important than feature $v$. To allow for controlled flexibility and avoid overly rigid constraints, we formalize this as the requirement that the expected difference in importance scores lies within a specified interval:
$\mathop{\mathbb{E}}_{\mathbf{X}\sim\mathcal{D}}\left[H_{v}(\mathbf{X},\theta)-H_{u}(\mathbf{X},\theta)\right]\in[\epsilon_{u,v},\delta_{u,v}],$ (9)
where $0<\epsilon_{u,v}\leq\delta_{u,v}$ are user-defined hyperparameters. The lower bound $\epsilon_{u,v}$ enforces a minimum meaningful margin of importance, ensuring $u$ is demonstrably more important than $v$. The upper bound $\delta_{u,v}$ prevents the optimization from pushing the difference to unrealistic extremes, which could destabilize training or force the model into pathological parameter regions. This interval-based formulation offers more nuanced control than a simple inequality constraint.
III-C Interpretability Loss
To integrate the DAG constraint from Eq. (9) into the training objective, we construct a differentiable interpretability loss function $\mathcal{H}:\Theta\to\mathbb{R}_{\geq 0}$. Since the expectation in Eq. (9) is over the data distribution, we work with its Monte Carlo estimate computed on a mini-batch.
For a batch of data $B=\{\mathbf{X}^{(1)},\dots,\mathbf{X}^{(|B|)}\}$ sampled i.i.d. from $\mathcal{D}$, we first compute the empirical mean importance difference for each edge $(u,v)\in E$:
$d_{u,v}(\theta;B)=\frac{1}{|B|}\sum_{\mathbf{X}\in B}\left[H_{v}(\mathbf{X},\theta)-H_{u}(\mathbf{X},\theta)\right].$ (10)
The batch interpretability loss is then defined as the sum of hinge-like penalties that activate when $d_{u,v}$ falls outside the desired interval $\epsilon_{u,v},\delta_{u,v}]$:
$\begin{split}\mathcal{H}(\theta;B)=\sum_{(u,v)\in E}\Big[&\max(\epsilon_{u,v}-d_{u,v}(\theta;B),\;0)\\ +&\max(d_{u,v}(\theta;B)-\delta_{u,v},\;0)\Big].\end{split}$ (11)
The first term penalizes differences that are too small (below $\epsilon_{u,v}$), and the second term penalizes differences that are too large (exceeding $\delta_{u,v}$). If $d_{u,v}(\theta;B)\in[\epsilon_{u,v},\delta_{u,v}]$, both terms evaluate to zero for that edge, incurring no penalty. The full loss $\mathcal{H}(\theta)$ is defined as the expectation of $\mathcal{H}(\theta;B)$ over batches $B\sim\mathcal{D}^{N}$. Crucially, because $H_{k}(\mathbf{X},\theta)$ is differentiable with respect to $\theta$ (flowing through $F_{\theta}$, the path integral, and the aggregation), the gradient $\nabla_{\theta}\mathcal{H}(\theta;B)$ can be computed efficiently via standard backpropagation through the entire computation graph.
III-D Gradient-Based Optimization via Projection Mapping
The IGBO framework gives rise to two distinct bi-objective optimization problems: (1) training the main model $F_{\theta}$ by minimizing $\mathcal{L}(\theta)$ and $\mathcal{H}(\theta)$ simultaneously, and (2) training the path generator $G(\theta_{G})$ by minimizing $\mathcal{L}_{\text{path}}$ and $\mathcal{L}_{\text{valid}}$. Both scenarios require combining gradients from competing objectives to find a parameter update direction that respects both.
Let $\mathbf{g}_{1}=\nabla_{\theta}\mathcal{J}_{1}$ and $\mathbf{g}_{2}=\nabla_{\theta}\mathcal{J}_{2}$ denote the gradients of two objectives $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ with respect to the parameters of interest (either $\theta$ or $\theta_{G}$).
Convergence Conditions: Before proceeding to the projection method, we first identify three conditions under which simultaneous descent is either impossible or unnecessary. If any of these conditions holds, the algorithm terminates early without proceeding to the projection step:
  1. 1.
    Maximally Conflicting Gradients: $\mathbf{g}_{1}\cdot\mathbf{g}_{2}=-\|\mathbf{g}_{1}\|\|\mathbf{g}_{2}\|$ (gradients are exactly opposite)
  2. 2.
    Negligible First Gradient: $\|\mathbf{g}_{1}\|<\epsilon$ (first objective is near optimum)
  3. 3.
    Negligible Second Gradient: $\|\mathbf{g}_{2}\|<\epsilon$ (second objective is near optimum)
In these cases, specialized handling is required rather than the projection method described below.
Fundamental Limitation of Convex Combinations: A naive approach uses a strict convex combination (in order to avoid checking many other cases $\lambda=0,1$ is ignored):
$\mathbf{v}_{\text{naive}}(\lambda)=\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2},\quad\lambda\in(0,1).$ (12)
However, this fails when gradients conflict: there exists no $\lambda\in(0,1)$ that guarantees $\mathbf{v}_{\text{naive}}(\lambda)\cdot\mathbf{g}_{1}>0$ and $\mathbf{v}_{\text{naive}}(\lambda)\cdot\mathbf{g}_{2}>0$ simultaneously. In fact:
  • For $\mathbf{g}_{1}\cdot\mathbf{g}_{2}\geq 0$: for every $\lambda\in(0,1)$ work
  • For $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$: just some of $\lambda\in(0,1)$ guarantees both
Complete Solution via Projection Mapping: We introduce a projection function $\mathcal{P}:\mathbb{R}^{p}\times\mathbb{R}^{p}\times(0,1)\to\mathbb{R}^{p}$ that maps any trade-off parameter $\lambda$ to a guaranteed descent direction:
$\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)=\begin{cases}\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2},\quad 0\leq g_{1}\cdot g_{2},\lambda\in(0,1)\\ \lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2},g_{1}\cdot g_{2}<0,\lambda\in(0,1)\end{cases}$ (13)
where:
$\mathbf{g}_{2}^{\perp 1}$ $=\mathbf{g}_{2}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{1}\|^{2}}\mathbf{g}_{1},$
$\mathbf{g}_{1}^{\perp 2}$ $=\mathbf{g}_{1}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{2}\|^{2}}\mathbf{g}_{2}.$
Geometric Intuition: The function $\mathcal{P}$ provides a complete parameterization of the simultaneous descent convex combinations $C_{\text{SD}}=\{\mathbf{v}=\alpha g_{1}+\beta g_{2}:\alpha,\beta>0,\,\mathbf{v}\cdot\mathbf{g}_{1}>0\text{ and }\mathbf{v}\cdot\mathbf{g}_{2}>0\}$:
  • When $\mathbf{g}_{1}\cdot\mathbf{g}_{2}>0$: $C_{\text{SD}}$ contains all convex combinations $\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2}$
  • When $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$: $C_{\text{SD}}$ contains $\lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2}$
  • When $\mathbf{g}_{1}\cdot\mathbf{g}_{2}=0$: $C_{\text{SD}}$ contains $\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2}$ for $\lambda\in(0,1)$
Unified Theoretical Guarantee (Complete Characterization): From Theorems A-A3 and A-A4:
$\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\in C_{\text{SD}}\quad\text{for all valid }\lambda$
and conversely, every $\mathbf{v}\in C_{\text{SD}}$ can be expressed as $\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)$ for some $\lambda$ and $\eta>0$.
Practical Implementation: The parameter $\lambda$ offers continuous trade-off control:
  • Fixed $\lambda$: Consistent balance (e.g., $\lambda=0.5$)
  • Scheduled $\lambda$: Adaptive balance during training
  • Dynamic $\lambda$: Responsive to gradient magnitudes or validation performance
Convergence Guarantees: Using Theorems A-A3 and A-A4, we establish:
  1. 1.
    Descent property: For any valid $\lambda$ and $\eta>0$:
    $\mathcal{J}_{1}(\theta-\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda))$ $=\mathcal{J}_{1}(\theta)-\eta C_{1}(\lambda)+O(\eta^{2})$
    $\mathcal{J}_{2}(\theta-\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda))$ $=\mathcal{J}_{2}(\theta)-\eta C_{2}(\lambda)+O(\eta^{2})$
    with $C_{1}(\lambda),C_{2}(\lambda)>0$.
  2. 2.
    Convergence: Any sequence $\{\lambda^{(k)}\}$ with valid values ensures convergence to Pareto-stationary points.
  3. 3.
    Robustness: The method tolerates gradient estimation errors (Theorem A-B1).
Algorithm 1 Projected Gradient Update via $\mathcal{P}$
1:Input: Gradients $\mathbf{g}_{1},\mathbf{g}_{2}$, trade-off $\lambda$, learning rate $\eta$
2:Output: Parameter update $\Delta\theta$
3:if $\mathbf{g}_{1}\cdot\mathbf{g}_{2}\geq 0$ then $\triangleright$ Aligned case ($\lambda\in(0,1)$)
4:  $\mathbf{v}\leftarrow\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2}$
[…]
6:  $\mathbf{g}_{2}^{\perp 1}\leftarrow\mathbf{g}_{2}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{1}\|^{2}}\mathbf{g}_{1}$
7:  $\mathbf{g}_{1}^{\perp 2}\leftarrow\mathbf{g}_{1}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{2}\|^{2}}\mathbf{g}_{2}$
8:  $\mathbf{v}\leftarrow\lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2}$
9:end if
[…]
11:Return $\Delta\theta$
This projection-based approach transforms the bi-objective optimization in IGBO from a heuristic balancing act to a principled procedure with full theoretical guarantees, enabling stable training with explicit trade-off control.
IV Theoretical Analysis
IV-A Computational Complexity Analysis
Training with IGBO introduces additional computational overhead compared to standard training, primarily due to the computation of Temporal Integrated Gradients (TIG) and the interpretability loss. We analyze the time and memory complexity for both standard training and IGBO.
Standard Training: For a sequential model processing input of length $T$, the per-iteration complexity is dominated by forward and backward propagation:
  • Forward Pass: $O(T\cdot F)$, where $F$ is the cost per time step.
  • Backward Pass (Primary Loss): $O(T\cdot F)$.
  • Total Time: $O(T\cdot F)$ per sample.
  • Memory: $O(T\cdot h)$ which $h$ is consumed additional memory per step.
IGBO Training: The additional computations in IGBO include:
  1. 1.
    Oracle Query: finding $K$ points which satisfies the mentioned conditions $O(K\cdot G)$ .
  2. 2.
    TIG Computation: Requires computing gradients of output w.r.t input at $M$ integration points. For one sample: $O(M\cdot T\cdot F)$.
  3. 3.
    Feature Importance $H_{k}$: Aggregation over time: $O(T^{2}\cdot d)$, where $d$ is feature dimension.
  4. 4.
    Interpretability Loss: Edge-wise comparisons: $O(|E|)$.
  5. 5.
    Gradient of $\mathcal{H}$: Backpropagation through TIG computation: $O(M\cdot T\cdot F)$.
Parallelization Effects:
  • Without GPU (Sequential):
    $\textbf{Time}=O(\text{MAX}\{M\cdot T\cdot F,G,T^{2}\cdot d,d^{2}\}),$
    $\text{ since }|E|\in O(d^{2})$
    (which is usually equivalent to $O(M\cdot T\cdot F+K\cdot G)$)
  • With GPU (Parallel):
    • The Oracle finding each point which satisfies the mentioned conditions: $G\rightarrow O(\mathcal{G})$.
    • TIG computations for different $\alpha$ values parallelizable: $M\rightarrow O(\log M)$
    • Feature Importance $H_{k}$ computations parallelizable: $T^{2}\cdot d\rightarrow O(\log T)$
    • Edge computations parallelizable: $|E|\rightarrow O(\log d)$
    • $O(\log d)\subset O(F)$ because in every forward step model cannot compute the scalar output in a time complexity better than $O(\log d)$.
    • Time: $O(\log M\cdot T\cdot F+K\cdot\mathcal{G})$
Memory Overhead: IGBO requires storing intermediate gradients for TIG computation and using the oracle:
  • Standard: $O(T\cdot h)$
  • IGBO (Sequential TIG computation + Oracle query): $O(T\cdot h+Q)$
  • IGBO (Parallel TIG computation + Oracle query): $O(M\cdot T\cdot h+Q)$ (for computing all $M$ points simultaneously) Additionally, with time complexity trade-off we have:
    • $\textbf{Memory}=O(M^{a}\cdot T\cdot h+Q)$
    • $\textbf{Time}=O(M^{1-a}\cdot T\cdot F+K\cdot\mathcal{G})$
    where $a\in(0,1)$.
Note on Geometric Update Complexity: The geometric update fucntion (Eqs. (13)) adds only $O(p)$ operations in sequential processing and $O(1)$ in parallel processing, where $p$ is the number of parameters. This is negligible compared to the $O(F)$, confirming that the geometric approach does not introduce significant computational burden beyond the interpretability loss computation.
V Experimental Design
This section outlines the comprehensive experimental plan to evaluate the IGBO framework. All experiments will be conducted following this protocol to ensure rigorous, reproducible validation.
V-A Research Objectives
Our experimental campaign aims to address four core objectives:
  • Objective 1: Constraint Enforcement Efficacy
    • Quantify how effectively IGBO enforces DAG-specified feature importance hierarchies
    • Measure the accuracy cost of enforcing interpretability constraints
  • Objective 2: Oracle Performance Analysis
    • Evaluate whether the Optimal Path Oracle generates more in-distribution integration paths compared to linear interpolation
    • Assess the impact of oracle-guided paths on attribution stability
  • Objective 3: Parameter Trade-off Characterization
    • Systematically analyze the effects of parameters $K$ (oracle points) and $M$ (total integration points)
    • Map the trade-off space between computational cost and path quality
  • Objective 4: Robustness Assessment
    • Test IGBO’s sensitivity to imperfect domain knowledge (noisy DAGs)
    • Evaluate robustness to hyperparameter variations ($\epsilon$, $\delta$, $\beta$)
V-B Dataset Preparation Plan
We will prepare and utilize three datasets:
  • Synthetic Hierarchical Dataset (SHD)
    • Generation Method: Create time-series data with known causal structure $X_{1}\rightarrow X_{2}\rightarrow X_{3}\rightarrow X_{4}$
    • Data Size: 10,000 sequences of length $T=50$ with $d=4$ features
    • Ground Truth: Precisely known feature importance hierarchy for validation
    • Split: 70% train, 15% validation, 15% test
  • MIMIC-III Clinical Dataset
    • Task: In-hospital mortality prediction within first 48 hours
    • Preprocessing: Extract 24-hour multivariate time-series, handle missing values with forward-fill, normalize features
    • DAG Construction: Collaborate with clinicians to define $G_{\text{clinical}}$: Vital signs $\succ$ Lab values $\succ$ Demographics
    • Expected Size: $\sim$15,000 patient stays after filtering
  • Financial Time-Series Dataset
    • Source: Yahoo Finance/Kaggle for S&P 500 constituents
    • Task: Next-day volatility classification (high/low)
    • Features: Technical indicators, price data, volume metrics
    • DAG: Define $G_{\text{finance}}$: Price-based $\succ$ Macroeconomic $\succ$ Sentiment
V-C Baseline Implementation Plan
We will implement and compare the following methods:
  • Primary Baselines:
    1. 1.
      Standard Model (STD): Base LSTM/Transformer trained only with task loss
    2. 2.
      Weighted Sum Regularization (WSR): $\mathcal{L}+\lambda\mathcal{H}$ with $\lambda\in\{0.001,0.01,0.1,1.0,10.0\}$
    3. 3.
      Post-hoc Fine-tuning (PHF): Two-stage: STD training followed by $\mathcal{H}$ fine-tuning
    4. 4.
      Linear Path IGBO (LP-IGBO): IGBO with $K=0$ (ablation study)
  • Implementation Details:
    • All methods use identical architecture: 2-layer LSTM, 128 hidden units, dropout=0.2
    • Adam optimizer, learning rate $5\times 10^{-4}$, batch size 32
    • Early stopping based on validation loss with patience=20
    • 5 random seeds per configuration for statistical significance
V-D Evaluation Metrics to Report
For each experiment, we will compute and report:
  • Primary Task Metrics:
    • Classification: Accuracy, F1-Score, AUROC (with confidence intervals)
    • Regression: RMSE, MAE, $R^{2}$
    • Relative performance: $\Delta_{\text{Acc}}=\frac{\text{Acc}_{\text{STD}}-\text{Acc}_{\text{Method}}}{\text{Acc}_{\text{STD}}}$
  • Interpretability Metrics:
    • DAG Satisfaction Rate (DSR): $\frac{1}{|E|}\sum\mathbb{I}(\hat{d}_{u,v}\in[\epsilon,\delta])$
    • Constraint Violation Magnitude (CVM): Average deviation from interval boundaries
    • Feature Importance Rank Correlation: Spearman correlation with ground truth (for SHD)
  • Oracle-Specific Metrics:
    • OOD Score: $\frac{1}{M}\sum_{j=1}^{M}D(\mathbf{q}_{j})$ where $D$ is pre-trained discriminator
    • Path Length Ratio: $\frac{\text{Oracle path length}}{\text{Linear path length}}$
    • Attribution Consistency: Std. dev. of $H_{k}$ scores across random seeds
    • Computational Cost: Training time per epoch, memory usage, inference time
V-E Parameter Studies Design
  • Study 1: Oracle Configuration ($K$, $M$)
    • Grid search: $K\in\{0,1,3,5,7\}$, $M\in\{10,20,50\}$
    • For each $(K,M)$: Train IGBO to convergence, record all metrics
    • Analyze: DSR vs. Computation time, OOD Score vs. $K$
  • Study 2: DAG Margin Sensitivity
    • Vary $\epsilon\in\{0.05,0.1,0.2\}$, $\delta\in\{0.3,0.5,1.0\}$
    • Fixed $K=3$, $M=20$
    • Measure: DSR, CVM, and accuracy trade-offs
  • Study 3: Robustness to Noisy Priors
    • Corruption levels: 0%, 10%, 20%, 30% of edges randomly reversed
    • Evaluate on both corrupted and uncorrupted edges separately
    • Assess recovery of correct constraints despite noise
  • Study 4: Generator Weight ($\beta$)
    • $\beta\in\{0.1,0.5,1.0,2.0,5.0\}$ in $\mathcal{L}_{G}=\mathcal{L}_{\text{path}}+\beta\mathcal{L}_{\text{valid}}$
    • Impact on path quality and training stability
V-F Experimental Procedure
  • Phase 0: Infrastructure Setup
    1. 1.
      Implement all models in PyTorch with modular design
    2. 2.
      Create data loaders for all three datasets
    3. 3.
      Implement metrics computation and logging utilities
    4. 4.
      Set up experiment tracking (Weights & Biases)
  • Phase 1: Baseline Establishment
    1. 1.
      Train Standard model on all datasets (5 seeds each)
    2. 2.
      Train WSR with all $\lambda$ values, identify Pareto frontier
    3. 3.
      Train PHF with different fine-tuning durations
    4. 4.
      Establish performance baselines
  • Phase 2: IGBO Core Evaluation
    1. 1.
      Pre-train discriminator $D$ for each dataset (50 epochs)
    2. 2.
      For each $(K,M)$ configuration in Study 1:
      • Train oracle generator $G$ (20 epochs warm-up)
      • Train full IGBO model
      • Record all metrics at convergence
    3. 3.
      Repeat for 5 random seeds per configuration
  • Phase 3: Sensitivity Analysis
    1. 1.
      Execute Study 2 (margin sensitivity)
    2. 2.
      Execute Study 3 (noisy DAG robustness)
    3. 3.
      Execute Study 4 (generator weight)
  • Phase 4: Analysis and Visualization
    1. 1.
      Statistical analysis: paired t-tests with Bonferroni correction
    2. 2.
      Generate all planned visualizations:
      • Pareto frontier plots (accuracy vs. DSR)
      • Trade-off surfaces ($K$, $M$, computation, DSR)
      • Feature importance heatmaps for case studies
      • t-SNE visualizations of integration paths
    3. 3.
      Compile results into comprehensive tables
V-G Validation Criteria
We will consider IGBO successful if:
  • Primary Success Criteria:
    • Achieves DSR ¿ 80% on test sets while maintaining ¡ 5% relative accuracy drop vs. STD
    • Oracle paths show significantly higher OOD Score than linear paths ($p<0.01$)
    • Attribution Consistency improves by ¿ 20% vs. linear paths
  • Secondary Success Criteria:
    • Parameters $K$ and $M$ show predictable trade-offs
    • Robust to moderate DAG noise (¡ 20% corruption)
    • Computationally feasible (training time ¡ 2× STD)
V-H Computational Resources and Timeline
  • Hardware:
    • 4× NVIDIA V100 GPUs (32GB memory each)
    • 64 CPU cores, 256GB RAM
    • Total estimated GPU hours: 2,000 hours
  • Software:
    • PyTorch 2.0, Python 3.9
    • Weights & Biases for experiment tracking
    • Docker containers for reproducibility
  • Timeline:
    • Months 1-2: Infrastructure and baseline implementation
    • Months 3-4: Core IGBO experiments (Phase 2)
    • Months 5: Sensitivity analyses (Phase 3)
    • Month 6: Analysis, visualization, and paper writing
V-I Reproducibility Measures
To ensure reproducibility:
  • All code will be publicly released on GitHub
  • Docker images with full environment specifications
  • Configuration files for every experiment
  • Pre-trained models and evaluation scripts
  • Detailed lab notebook documenting all experimental decisions
This experimental design provides a complete roadmap for evaluating IGBO. The structured approach ensures comprehensive assessment of all framework components while maintaining scientific rigor and reproducibility.
VI Conclusion and Future Work
We have introduced Interpretability-Guided Bi-objective Optimization (IGBO), a novel framework for training machine learning models to adhere to domain-specific interpretability constraints. By formalizing these constraints via a Directed Acyclic Graph (DAG) acting on a differentiable feature importance function $H_{k}(\mathbf{X},\theta)$, and by deriving a geometrically motivated update rule that dynamically decomposes task and interpretability gradients, IGBO provides a principled method to navigate the accuracy-interpretability trade-off. Theoretical analysis guarantees convergence to stationary points with clear semantic meanings: joint optima, task-only optima, or Pareto-optimal trade-off points.
The empirical framework we have outlined is designed to validate IGBO’s effectiveness in enforcing structured priors with minimal accuracy loss, outperforming naive regularization baselines. The ability to explicitly encode relational domain knowledge (“feature A should be more important than feature B”) and have the learning algorithm respect it during training represents a significant step towards more transparent, trustworthy, and alignable models for high-stakes applications.
Limitations and Future Work:
  • Computational Overhead: The computation of $H_{k}(\mathbf{X},\theta)$, involving multiple gradient evaluations for the path integral, adds overhead per training step. While parallelizable, future work should investigate more efficient, approximate attribution methods that retain differentiability but are cheaper to compute.
  • DAG Specification: The current framework requires the DAG and intervals $\epsilon,\delta]$ to be specified by an expert. Future work could explore learning or refining this graph from data, or from interactions with a human-in-the-loop.
  • Extension to More Objectives: The geometric update rule is formulated for two objectives. Extending this principle to handle three or more concurrent objectives (e.g., accuracy, interpretability, fairness, robustness) is a challenging but promising direction for building models that satisfy multiple desiderata simultaneously.
  • Theoretical Refinements: Further analysis could establish non-asymptotic convergence rates or study the geometry of the Pareto frontier discovered by IGBO.
In conclusion, IGBO offers a rigorous and practical pathway to move beyond post-hoc explanations, enabling the creation of models whose internal decision-making processes are inherently guided by, and thus more aligned with, human-understandable semantic structures.
Appendix A Proofs
This appendix contains the complete proofs for the theoretical claims made in Section III-D.
A-A Proofs for Projection Mapping Properties
A-A1 Preliminary Definitions
Let $\mathbf{g}_{1}=\nabla_{\theta}\mathcal{J}_{1}$ and $\mathbf{g}_{2}=\nabla_{\theta}\mathcal{J}_{2}$ be two non-zero gradient vectors. Define the set of simultaneous descent directions as:
$C_{\text{SD}}=\{\mathbf{v}=\alpha\mathbf{g}_{1}+\beta\mathbf{g}_{2}:\alpha,\beta>0,\,\mathbf{v}\cdot\mathbf{g}_{1}>0\text{ and }\mathbf{v}\cdot\mathbf{g}_{2}>0\}.$ (14)
A-A2 Lemma 1: Limitation of Naive Convex Combination
Claim: When $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$, there exists no $\lambda\in(0,1)$ such that the naive convex combination $\mathbf{v}_{\text{naive}}(\lambda)=\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2}$ guarantees $\mathbf{v}_{\text{naive}}(\lambda)\cdot\mathbf{g}_{1}>0$ and $\mathbf{v}_{\text{naive}}(\lambda)\cdot\mathbf{g}_{2}>0$ simultaneously for all $\lambda\in(0,1)$. Only some values of $\lambda$ satisfy both conditions.
Proof.
Compute the dot products:
$\mathbf{v}_{\text{naive}}(\lambda)\cdot\mathbf{g}_{1}$ $=\lambda\|\mathbf{g}_{1}\|^{2}+(1-\lambda)(\mathbf{g}_{1}\cdot\mathbf{g}_{2}),$
$\mathbf{v}_{\text{naive}}(\lambda)\cdot\mathbf{g}_{2}$ $=\lambda(\mathbf{g}_{1}\cdot\mathbf{g}_{2})+(1-\lambda)\|\mathbf{g}_{2}\|^{2}.$
For $\mathbf{v}_{\text{naive}}(\lambda)$ to be a simultaneous descent direction, we need both expressions positive. This gives two inequalities:
1. $\lambda\|\mathbf{g}_{1}\|^{2}+(1-\lambda)(\mathbf{g}_{1}\cdot\mathbf{g}_{2})>0$ 2. $\lambda(\mathbf{g}_{1}\cdot\mathbf{g}_{2})+(1-\lambda)\|\mathbf{g}_{2}\|^{2}>0$
Since $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$, the first inequality implies:
$\lambda>\frac{-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}{\|\mathbf{g}_{1}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}.$
The second inequality implies:
$\lambda<\frac{\|\mathbf{g}_{2}\|^{2}}{\|\mathbf{g}_{2}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}.$
Thus, $\lambda$ must satisfy:
$\frac{-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}{\|\mathbf{g}_{1}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}<\lambda<\frac{\|\mathbf{g}_{2}\|^{2}}{\|\mathbf{g}_{2}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}.$
This interval is non-empty only if:
$\frac{-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}{\|\mathbf{g}_{1}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}<\frac{\|\mathbf{g}_{2}\|^{2}}{\|\mathbf{g}_{2}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}.$
Cross-multiplying (note denominators are positive since $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$) yields:
$(\mathbf{g}_{1}\cdot\mathbf{g}_{2})(\|\mathbf{g}_{2}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2}))<\|\mathbf{g}_{2}\|^{2}(\|\mathbf{g}_{1}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})).$
Simplifying:
$(\mathbf{g}_{1}\cdot\mathbf{g}_{2})\|\mathbf{g}_{2}\|^{2}+(\mathbf{g}_{1}\cdot\mathbf{g}_{2})^{2}<\|\mathbf{g}_{1}\|^{2}\|\mathbf{g}_{2}\|^{2}-\|\mathbf{g}_{2}\|^{2}(\mathbf{g}_{1}\cdot\mathbf{g}_{2}).$
Cancelling $(\mathbf{g}_{1}\cdot\mathbf{g}_{2})\|\mathbf{g}_{2}\|^{2}$ on both sides gives:
$(\mathbf{g}_{1}\cdot\mathbf{g}_{2})^{2}<\|\mathbf{g}_{1}\|^{2}\|\mathbf{g}_{2}\|^{2},$
which always holds by the Cauchy-Schwarz inequality when $\mathbf{g}_{1}$ and $\mathbf{g}_{2}$ are not collinear. However, when $\mathbf{g}_{1}$ and $\mathbf{g}_{2}$ are exactly opposite ($\mathbf{g}_{1}\cdot\mathbf{g}_{2}=-\|\mathbf{g}_{1}\|\|\mathbf{g}_{2}\|$), the interval collapses. For general conflicting gradients ($\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$), the interval exists but does not cover all $\lambda\in(0,1)$. Thus, not every $\lambda\in(0,1)$ yields a simultaneous descent direction. ∎
A-A3 Theorem 1: Aligned Gradients Case
Claim: If $\mathbf{g}_{1}\cdot\mathbf{g}_{2}\geq 0$, then for any $\lambda\in(0,1)$, the vector $\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)=\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2}$ belongs to $C_{\text{SD}}$. Moreover, every vector in $C_{\text{SD}}$ can be expressed as $\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)$ for some $\lambda\in(0,1)$ and $\eta>0$.
Proof.
First, verify the descent property. For $\mathbf{v}=\lambda\mathbf{g}_{1}+(1-\lambda)\mathbf{g}_{2}$:
$\mathbf{v}\cdot\mathbf{g}_{1}$ $=\lambda\|\mathbf{g}_{1}\|^{2}+(1-\lambda)(\mathbf{g}_{1}\cdot\mathbf{g}_{2})\geq\lambda\|\mathbf{g}_{1}\|^{2}>0,$
$\mathbf{v}\cdot\mathbf{g}_{2}$ $=\lambda(\mathbf{g}_{1}\cdot\mathbf{g}_{2})+(1-\lambda)\|\mathbf{g}_{2}\|^{2}\geq(1-\lambda)\|\mathbf{g}_{2}\|^{2}>0,$
since $\lambda\in(0,1)$, $\|\mathbf{g}_{1}\|,\|\mathbf{g}_{2}\|>0$, and $\mathbf{g}_{1}\cdot\mathbf{g}_{2}\geq 0$. Hence $\mathbf{v}\in C_{\text{SD}}$.
For the converse, let $\mathbf{w}=\alpha\mathbf{g}_{1}+\beta\mathbf{g}_{2}\in C_{\text{SD}}$ with $\alpha,\beta>0$. Choose $\lambda=\frac{\alpha}{\alpha+\beta}\in(0,1)$ and $\eta=\alpha+\beta>0$. Then:
$\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)=(\alpha+\beta)\left(\frac{\alpha\mathbf{g}_{1}+\beta\mathbf{g}_{2}}{\alpha+\beta}\right)=\alpha\mathbf{g}_{1}+\beta\mathbf{g}_{2}=\mathbf{w}.$
Thus $\mathbf{w}$ has the desired form. ∎
A-A4 Theorem 2: Conflicting Gradients Case
Claim: If $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$, then for any $\lambda\in(0,1)$, the vector
$\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)=\lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2},$
where $\mathbf{g}_{2}^{\perp 1}=\mathbf{g}_{2}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{1}\|^{2}}\mathbf{g}_{1}$ and $\mathbf{g}_{1}^{\perp 2}=\mathbf{g}_{1}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{2}\|^{2}}\mathbf{g}_{2}$, belongs to $C_{\text{SD}}$. Moreover, every vector in $C_{\text{SD}}$ can be expressed as $\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)$ for some $\lambda\in(0,1)$ and $\eta>0$.
Proof.
First, note that $\mathbf{g}_{2}^{\perp 1}$ is orthogonal to $\mathbf{g}_{1}$, and $\mathbf{g}_{1}^{\perp 2}$ is orthogonal to $\mathbf{g}_{2}$. Compute:
$\mathbf{g}_{2}^{\perp 1}\cdot\mathbf{g}_{1}$ $=0,$
$\mathbf{g}_{1}^{\perp 2}\cdot\mathbf{g}_{2}$ $=0.$
Now let $\mathbf{v}=\lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2}$. Then:
$\mathbf{v}\cdot\mathbf{g}_{1}$ $=\lambda(\mathbf{g}_{2}^{\perp 1}\cdot\mathbf{g}_{1})+(1-\lambda)(\mathbf{g}_{1}^{\perp 2}\cdot\mathbf{g}_{1})=(1-\lambda)\|\mathbf{g}_{1}^{\perp 2}\|^{2},$
$\mathbf{v}\cdot\mathbf{g}_{2}$ $=\lambda(\mathbf{g}_{2}^{\perp 1}\cdot\mathbf{g}_{2})+(1-\lambda)(\mathbf{g}_{1}^{\perp 2}\cdot\mathbf{g}_{2})=\lambda\|\mathbf{g}_{2}^{\perp 1}\|^{2}.$
Since $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$, we have $\mathbf{g}_{1}^{\perp 2}\neq\mathbf{0}$ and $\mathbf{g}_{2}^{\perp 1}\neq\mathbf{0}$ (unless $\mathbf{g}_{1}$ and $\mathbf{g}_{2}$ are collinear, which is excluded by the condition $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$ and not exactly opposite per termination condition 1). Hence $\|\mathbf{g}_{1}^{\perp 2}\|^{2}>0$ and $\|\mathbf{g}_{2}^{\perp 1}\|^{2}>0$. For $\lambda\in(0,1)$, both dot products are positive, so $\mathbf{v}\in C_{\text{SD}}$.
For the converse, let $\mathbf{w}=\alpha\mathbf{g}_{1}+\beta\mathbf{g}_{2}\in C_{\text{SD}}$ with $\alpha,\beta>0$. We seek $\lambda\in(0,1)$ and $\eta>0$ such that:
$\eta\left(\lambda\mathbf{g}_{2}^{\perp 1}+(1-\lambda)\mathbf{g}_{1}^{\perp 2}\right)=\alpha\mathbf{g}_{1}+\beta\mathbf{g}_{2}.$
Express $\mathbf{g}_{1}^{\perp 2}$ and $\mathbf{g}_{2}^{\perp 1}$ in terms of $\mathbf{g}_{1}$ and $\mathbf{g}_{2}$:
$\mathbf{g}_{1}^{\perp 2}$ $=\mathbf{g}_{1}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{2}\|^{2}}\mathbf{g}_{2},$
$\mathbf{g}_{2}^{\perp 1}$ $=\mathbf{g}_{2}-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{1}\|^{2}}\mathbf{g}_{1}.$
Substitute into the equation and equate coefficients of $\mathbf{g}_{1}$ and $\mathbf{g}_{2}$:
$\eta\left(\lambda\left(-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{1}\|^{2}}\right)+(1-\lambda)\right)$ $=\alpha,$
$\eta\left(\lambda+(1-\lambda)\left(-\frac{\mathbf{g}_{1}\cdot\mathbf{g}_{2}}{\|\mathbf{g}_{2}\|^{2}}\right)\right)$ $=\beta.$
This is a linear system in $\eta$ and $\lambda$. Solving yields:
$\lambda=\frac{\alpha\|\mathbf{g}_{1}\|^{2}\|\mathbf{g}_{2}\|^{2}-\alpha\|\mathbf{g}_{2}\|^{2}(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}{(\alpha+\beta)\|\mathbf{g}_{1}\|^{2}\|\mathbf{g}_{2}\|^{2}-(\mathbf{g}_{1}\cdot\mathbf{g}_{2})(\alpha\|\mathbf{g}_{2}\|^{2}+\beta\|\mathbf{g}_{1}\|^{2})},$
$\eta=\frac{\alpha\|\mathbf{g}_{1}\|^{2}-\beta(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}{\|\mathbf{g}_{1}\|^{2}-\lambda(\mathbf{g}_{1}\cdot\mathbf{g}_{2})}.$
Since $\alpha,\beta>0$ and $\mathbf{g}_{1}\cdot\mathbf{g}_{2}<0$, the denominator for $\lambda$ is positive, and $\lambda\in(0,1)$. Also $\eta>0$. Thus every $\mathbf{w}\in C_{\text{SD}}$ can be represented as $\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)$. ∎
A-A5 Corollary 1: Descent Property
Claim: For any valid $\lambda\in(0,1)$ and step size $\eta>0$ sufficiently small, the update $\theta^{\prime}=\theta-\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)$ decreases both objectives:
$\mathcal{J}_{1}(\theta^{\prime})$ $=\mathcal{J}_{1}(\theta)-\eta C_{1}(\lambda)+O(\eta^{2}),$
$\mathcal{J}_{2}(\theta^{\prime})$ $=\mathcal{J}_{2}(\theta)-\eta C_{2}(\lambda)+O(\eta^{2}),$
where $C_{1}(\lambda),C_{2}(\lambda)>0$.
Proof.
By Taylor expansion:
$\mathcal{J}_{1}(\theta^{\prime})=\mathcal{J}_{1}(\theta)-\eta\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{1}+O(\eta^{2}).$
From Theorems A-A3 and A-A4, we have $\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{1}>0$. Set $C_{1}(\lambda)=\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{1}>0$. Similarly for $\mathcal{J}_{2}$ with $C_{2}(\lambda)=\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)\cdot\mathbf{g}_{2}>0$. ∎
A-B Proofs for Noise Robustness Properties
A-B1 Theorem 3: Robustness to Gradient Noise
Claim: Let $\tilde{\mathbf{g}}_{1}=\mathbf{g}_{1}+\boldsymbol{\epsilon}_{1}$, $\tilde{\mathbf{g}}_{2}=\mathbf{g}_{2}+\boldsymbol{\epsilon}_{2}$ be noisy gradient estimates with $\mathbb{E}[\boldsymbol{\epsilon}_{i}]=0$ and $Var(\boldsymbol{\epsilon}_{i})=\sigma_{i}^{2}I$. Then the expected update direction $\mathbb{E}[\mathcal{P}(\tilde{\mathbf{g}}_{1},\tilde{\mathbf{g}}_{2},\lambda)]$ is $\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)+O(\sigma^{2})$, and the variance is $O(\sigma^{2})$.
Proof.
The projection function $\mathcal{P}$ is smooth (piecewise linear) in its arguments. By Taylor expansion around $(\mathbf{g}_{1},\mathbf{g}_{2})$:
$\mathcal{P}(\tilde{\mathbf{g}}_{1},\tilde{\mathbf{g}}_{2},\lambda)=\mathcal{P}(\mathbf{g}_{1},\mathbf{g}_{2},\lambda)+\sum_{i=1}^{2}\frac{\partial\mathcal{P}}{\partial\mathbf{g}_{i}}\boldsymbol{\epsilon}_{i}+O(\|\boldsymbol{\epsilon}\|^{2}).$
Taking expectation, the linear terms vanish since $\mathbb{E}[\boldsymbol{\epsilon}_{i}]=0$, leaving bias of order $\sigma^{2}$. The variance comes from the linear terms: $Var\left(\frac{\partial\mathcal{P}}{\partial\mathbf{g}_{i}}\boldsymbol{\epsilon}_{i}\right)=O(\sigma_{i}^{2})$. ∎
A-B2 Corollary 2: Effect of Batch Size
Claim: If gradients are estimated using mini-batches of size $B$, then $\sigma^{2}\propto 1/B$. Thus both bias and variance of the update direction scale as $O(1/B)$.
Proof.
For i.i.d. samples, the variance of the gradient estimate scales inversely with batch size: $\sigma^{2}=\sigma_{0}^{2}/B$. Substituting into Theorem A-B1 gives the result. ∎